### v1

# AWARE: Audio Watermarking via Adversarial Resistance to Edits

Abstract
Prevailing practice in learning-based audio watermarking is to pursue robustness by expanding the set of simulated distortions during training. However, such surrogates are narrow and prone to overfitting. This paper presents AWARE (Audio Watermarking with Adversarial Resistance to Edits), an alternative approach that avoids reliance on attack-simulation stacks and handcrafted differentiable distortions. Embedding is obtained via adversarial optimization in the time–frequency domain under a level-proportional perceptual budget. Detection employs a time–order–agnostic detector with a Bitwise Readout Head (BRH) that aggregates temporal evidence into one score per watermark bit, enabling reliable watermark decoding even under desynchronization and temporal cuts. Empirically, AWARE attains high audio quality and speech intelligibility (PESQ/STOI) and consistently low BER across various audio edits, often surpassing representative state-of-the-art learning-based audio watermarking systems.
1 Introduction
Digital watermarking experienced its first major wave of research activity in the 1990s alongside the rapid proliferation of the Internet. Early systems were primarily designed for copyright protection and digital rights management (DRM), with the seminal work of Cox et al. introducing spread-spectrum principles to watermarking and setting the agenda for robustness-focused design [1]. Subsequent developments broadened the methodological toolbox with techniques such as quantization index modulation (QIM) [2] and patchwork-style techniques [3]. While initial approaches were conceived as modality-agnostic and applicable across multimedia, the field soon bifurcated into image- and audio-specific lines of work [4, 5], each exploiting modality characteristics to improve embedding efficiency and detection reliability. The dominant use case remained copyright protection, driven by the rise of large-scale online content distribution and associated piracy.
Despite their impact, traditional watermarking techniques (particularly in audio) face persistent limitations. Beyond limited robustness to classical signal processing, two challenges loom large: desynchronization and waveform cuts. Systems often include dedicated synchronization codes to address time-scale modifications, resampling drift, cropping, and jitter. However, reliably detecting synchronization markers is nearly as difficult as extracting the watermark itself, and thus inherits similar failure modes under distortion. Moreover, many legacy designs embed watermark bits within a single frame or a narrow group of frames, yielding limited temporal redundancy and weak fragment-level detectability. Missing or re-ordered frames, or partial content removal, can therefore break the decoding process and make the watermark hard or impossible to reconstruct.
The advent of modern generative AI has precipitated a renaissance in digital audio watermarking. High-fidelity synthesis models such as: GANs [6] and diffusion models [7] enable convincing audio and audiovisual “deepfakes” at scale. The risks span reputational harm, fraud, misinformation, and weakened evidence reliability. In response, watermarking has re-emerged as a practical mechanism to label both synthetic and authentic content to support provenance, traceability, and downstream moderation. Accordingly, policy frameworks increasingly cite watermarking among key techniques for AI transparency and content provenance [8].
Concurrently, the community has begun to “fight fire with fire,” developing end-to-end deep learning (DL) audio watermarking systems. Leading this line of research, RobustDNN [9] defined the basic blueprint, after which WavMark [10] and AudioSeal [11] introduced meaningful improvements. Nevertheless, contemporary benchmark studies still indicate unresolved limitations, with a clear room for progress [12, 13].
Critically, watermark decoding for audio under temporal cuts remains underexplored, aside from zero-bit approaches [14] that have limited practical scope. Unlike images, where spatial cropping typically retains substantial context, audio pipelines frequently yield spliced content: selective cuts, concatenation of short segments from different sources, or montage-like edits. Such edits can sound natural to human listeners, yet disrupt global synchronization and erase large portions of the embedded message. Practical deployments therefore require watermarking that survives cuts and splicing. Meanwhile, high-fidelity voice cloning introduces a new problem that challenges established provenance mechanisms.
A general trend in deep learning has been to port high-performing architectures from other domains, most notably computer vision, by “stacking layers” from CNN/Transformer backbones. This pattern has also influenced DL-based audio watermarking. Classic signal-processing–driven designs, by contrast, were intentionally constrained: each design choice (domain selection, embedding operator, synchronization mechanism, detection statistic) was motivated by the signal model and anticipated threat model.
While modern DL-based approaches have accelerated progress, they often overlook the unique nature of watermark detection in audio. Watermark detection differs fundamentally from object/keyword detection or semantic classification. The target is not a semantic entity localized in space or time. It is a weak, distributed pattern, encoding bits that must be sequence-consistent under time-warping and cutting. The decoder must aggregate evidence over time, maintain or recover alignment, and ultimately produce a bitstream (with reliability scores).
To address these gaps, this paper introduces AWARE (Audio Watermarking with Adversarial Resistance to Edits): an adversarial watermarking procedure "aware" of auditory perception and audio signal structure. It employs an adversarial embedding procedure under a level-proportional perceptual budget and a time–order–agnostic detector with a Bitwise Readout Head that aggregates temporal evidence into per-bit scores. Comprehensive experiments show high perceptual quality and intelligibility with consistently low BER under diverse edits, often surpassing state-of-the-art baselines and yielding a more stable BER profile across conditions, including desynchronization and temporal cuts.
2 Background
In the standard adversarial setting, we start with a clean input $x$, its label $y$, a model $f$, and a task-specific loss function $\mathcal{L}$. The objective is to find a small, norm-bounded perturbation $\Delta$ that maximizes the loss $\mathcal{L}$:
$\max_{\lVert\Delta\rVert_{p}\leq\epsilon}\ \mathcal{L}\big(f(x+\Delta),\,y\big).$ (1)
Intuitively, we seek the smallest change (constrained by $\|\Delta\|_{p}\leq\varepsilon$) that maximizes the degradation of the model’s performance on $x$. The choice of $p$ (e.g., $p=\infty$ or $2$) and radius $\varepsilon$ encodes the allowable perturbation “budget”.
Adversarial perturbations have proved effective across diverse tasks within the audio modality [15, 16, 17], predominantly as tools to break models rather than to enforce desired behaviors. Li et al. [18] argue that watermarking is essentially a signal perturbation optimized to steer the detector’s outputs. Hence, they introduce adversarial shallow watermarking for images, pairing a frozen detector with an embedding procedure that adjusts the carrier until a randomly initialized detector yields the desired watermark bits. Two insights are particularly relevant to our setting: (i) detector architectures should remain shallow to improve out-of-distribution generalization and resilience to unseen attacks, and (ii) distortion-simulation stacks can be simplified or eliminated, avoiding reliance on large suites of handcrafted differentiable attacks.
Learning-based watermarking often gains robustness by inserting differentiable “attack layers” into training. However, this strategy faces practical limits: the space of plausible attacks is vast (combinatorial in type and composition), so exhaustive coverage during training is infeasible. Moreover, Li et al. observe that DL-based systems tend to overfit to the attack set seen during training and generalize poorly when novel perturbations are introduced.
This adversarial approach has another advantage that it enables rapid evaluation of different architectural choices under a broad set of perturbations by avoiding lengthy supervised training. The resulting insights can later inform the design of learning-based watermarking systems (with or without attack layers).
Naively adapting the Li et al. approach to audio waveforms/spectrograms inherits the same issues seen in many DL-based audio watermarking systems under cuts and desynchronization, as the method was tailored to images. Audio watermark decoders require architectural mechanisms that are intrinsically robust to common edits and temporal misalignments. If the architecture is not inherently robust, no amount of augmentation or attack-layer engineering will make training reliably effective.
3 AWARE: Method
Watermark Embedding. Embedding is carried out in the time-frequency (TF) domain, a standard and effective setting for audio watermarking. The complete procedure is outlined in Algorithm 1.
Algorithm 1 AWARE Embedding Procedure
1:waveform $x$, watermark $\tilde{w}\in\{-1,+1\}^{N}$, detector $D$, embedding band $\mathcal{F}=[f_{\ell},f_{h}]$, bin budgets $B$, margin weight $\lambda$, iterations $K$.
[…]
4:for $k=1$ to $K$ do
5:  $y\leftarrow D(M+\Delta)$
6:  $\mathcal{L}\leftarrow\frac{1}{N}\|y-\tilde{w}\|_{2}^{2}-\lambda\frac{1}{N}\sum_{i}|y_{i}|$ $\triangleright$ push loss
7:  $\Delta\leftarrow\mathrm{OptimizerStep}(\Delta,\nabla_{\Delta}\mathcal{L})$
8:  for all $(f,u)$: $\Delta_{f,u}\leftarrow\mathrm{clip}(\Delta_{f,u},-B_{f,u},+B_{f,u})$
9:end for
[…]
12:watermarked audio $\tilde{x}$
Let $x\in\mathbb{R}^{T}$ be a waveform and let $\lvert\mathrm{STFT}(x)\rvert\in\mathbb{R}_{\geq 0}^{F\times U}$ denote its short-time Fourier transform (STFT) magnitude, with frequency bins $f\in[0,\dots,F-1]$ and time frames $u\in[0,\dots,U-1]$. We restrict perturbations to an audible midband $\mathcal{F}=\{f:f_{\ell}\leq f\leq f_{h}\}$ with $f_{\ell}=500~\text{Hz}$ and $f_{h}=4000~\text{Hz}$ to avoid removal by low/high-pass filters.
Phase-domain watermark embedding is avoided due to human hearing being largely insensitive to changes in phase, which allows for low-audibility phase manipulations that effectively erase the mark. Accordingly, we modify magnitudes under perceptual constraints and preserve the original phase for reconstruction (iSTFT).
We denote $w$ as a watermark of $N$ bits, and $D$ represents a randomly initialized detector with frozen weights that maps a TF magnitude representation of $x$ to $\left(-1,+1\right)^{N}$. Watermark bits are encoded antipodally as $\tilde{w}\in\{-1,+1\}^{N}$. This centers targets at zero, yields balanced gradients, and pairs naturally with margin-based objectives and sign decoding, unlike $\{0,1\}$ coding which pushes scores towards probability bounds and can hinder optimization.
The embedding procedure minimizes a push loss objective that drives the detector toward accurate and confident bipolar decisions on the target bits (increasing the margin to $\pm 1$). Let $M=\lvert\mathrm{STFT}(x)\rvert$ be the STFT magnitude, and let $\Delta$ be a magnitude perturbation supported on $\mathcal{F}$. The detector prediction is $y=D(M+\Delta)\in\left(-1,+1\right)^{N}$. Optimization objective is given by:
$\mathcal{L}_{\text{push}}(M,\Delta;\tilde{w})\;=\;\underbrace{\frac{1}{N}\|y-\tilde{w}\|_{2}^{2}}_{\text{MSE to targets}}\;-\;\lambda\,\underbrace{\frac{1}{N}\sum_{i=1}^{N}|y_{i}|}_{\text{margin term}},$ (2)
with $\lambda>0$ controlling the margin strength.
Rather than imposing a single global norm budget on $\Delta$, we use a per-bin, level-proportional budget that allows larger changes where the signal is louder and smaller changes in quiet regions, consistent with basic psychoacoustics. Let $\tau_{\text{dB}}>0$ be a tolerance parameter (in dB). Its linear amplitude factor is $\eta\;=\;10^{-\tau_{\text{dB}}/20}.$ For each TF bin $(f,u)$, the admissible magnitude change is bounded by $|\Delta_{f,u}|\;\leq\;\eta\,M_{f,u}$. We enforce these bounds via projection (clipping) after each optimizer step. Unlike loss-only quality terms (as in Li et al.), this explicit budget provides stronger and more reliable quality control and avoids re-embedding caused by overly aggressive updates. Bin-wise perceptual budgeting is common in classical, masking-inspired watermarking, but it has been largely de-emphasized in recent DL-based systems, where perceptual control is typically folded into a soft term in the overall loss that offers weaker guarantees than the hard constraints. In future work the tolerance $\tau_{\text{dB}}$ can be replaced by frequency-dependent thresholds or full psychoacoustic models (threshold-in-quiet and simultaneous masking), enabling budgets that also consider neighboring-frequency masking.
Figure 1: AWARE detector architecture.
Detector architecture. The detector architecture (illustrated in Figure  1) is designed so that activations across all layers remain stable under common audio edits. Robustness thus derives from the architecture itself rather than from optimization or distortion–simulation surrogates which tend to drive the system toward overfitting.
The first layer of the detector computes a Mel–spectrogram from the STFT magnitude. The Mel domain aggregates spectral energy into perceptually motivated bands, yielding a representation that is more robust than raw spectrograms to mild time–frequency distortions. Moreover, Mel features are standard in TTS/vocoder pipelines, increasing the chance that a watermark detectable in Mel-space survives voice cloning.
Following ablation findings in Li et al., we include a single pooling layer. This coarsens temporal resolution and improves robustness to local jitter and minor desynchronization.
The next stage contains three feature extraction blocks that intentionally avoid temporal mixing by treating Mel bands as channels and applying convolution along time dimension (kernel size =1=1, stride =1=1), followed by instance normalization and a Leaky ReLU activation. In this way, each time frame is processed independently, so the convolution kernels act as channel mixers across frequency at a fixed time, thereby avoiding sensitivity to cuts and re-ordering, that would otherwise induce activation-statistics drift.
Fully connected (FC) layers bind decisions to absolute positions and fixed input lengths. Deletions and splicing change the index–to–time mapping and lead to brittle, non-invariant activations. Moreover, FC layers require length standardization or padding that obscures real edit patterns. We therefore exclude FC layers entirely from the detector.
Instead, we introduce a Bitwise Readout Head (BRH) that reads out $N$ bits using paired convolutional filters. These filters aggregate evidence over time, and produce one position-agnostic score per bit. Concretely, the BRH applies two filter banks to the extracted features $Z\in\mathbb{R}^{C\times U^{\prime}}$:
$A^{(0)}$ $=W^{(0)}Z\in\mathbb{R}^{N\times U^{\prime}},$ (3)
$A^{(1)}$ $=W^{(1)}Z\in\mathbb{R}^{N\times U^{\prime}}$
with $W^{(0)},W^{(1)}\in\mathbb{R}^{N\times C}$, so that for each bit index $i$ we obtain two activation traces $A^{(0)}_{i,\cdot}$ and $A^{(1)}_{i,\cdot}$ along time. The traces are then aggregated by global averaging:
$\bar{a}^{(b)}_{i}\;=\;\frac{1}{U^{\prime}}\sum_{u=1}^{U^{\prime}}A^{(b)}_{i,u},\quad b\in\{0,1\},$ (4)
and contrasted to produce a single bit score:
$g_{i}\;=\;\bar{a}^{(1)}_{i}-\bar{a}^{(0)}_{i},\qquad y_{i}\;=\;\tanh(g_{i}).$ (5)
Here, $\tanh(\cdot)$ is a monotone squashing function that maps scores to (1,+1)(-1,+1). The output $y=(y_{1},\ldots,y_{N})$ thus provides one position-agnostic score per watermark bit, obtained by temporal evidence aggregation within the BRH.
Having described the mechanics of the BRH, we now sketch the intuition behind its structure. Convolutional filters are typically crafted to fire on specific stimuli or patterns (e.g., in image classification/object detection, some filters respond to “ears”, others to “eyes”, etc.). In watermark detection, those stimuli are bits. Accordingly, the BRH allocates two filters per bit $b$: one tuned to evidence for $b=1$ and one for $b=0$. Each filter produces a temporal activation trace that global averaging converts into evidence scores. The bit decision becomes a simple competition. Whichever filter accumulates more evidence over time, “wins”.
Global averaging in BRH removes the time axis before the final decision, ensuring that cropping, splicing, or frame deletions change the amount of evidence but not its required position. Because the feature extractor never relies on multi-frame receptive fields (kernel size for all convs is $1$), activation statistics remain stable under variable-length inputs and missing frames, thus enabling reliable fragment-level detection.
4 Experimental Setup
We compare against some of the strongest publicly available baselines (WavMark and AudioSeal) on VCTK [19] and LibriSpeech [20] datasets at $16$ kHz sampling rate. Perceptual quality is evaluated with PESQ [21] and STOI [22], while watermark robustness is measured by bit error rate (BER).
For a fair comparison, all methods are tested at a payload of $16$ bps, matching the training configuration of the comparative baselines, although AWARE exhibits similar performance at capacities exceeding $20$ bps. Adversarial embedding is optimized for $K=500$ iterations using the NAdam optimizer [23] (learning rate $0.1$), with a reduce-on-plateau scheduler (factor $0.9$). The push-loss margin weight is set to $\lambda=0.1$.
[…]
5 Results and Analysis
Table 1 shows that AWARE achieves high audio quality and speech intelligibility, although slightly below AudioSeal, but above WavMark. Nevertheless, these scores indicate that the watermark remains imperceptible or near-imperceptible at all times. Given our goal of stronger cross-attack robustness, a modest quality compromise is expected.
Table 1: Speech quality and intelligibility (mean $\pm$ std).
Method PESQ $\uparrow$ STOI $\uparrow$
WavMark $3.96\pm 0.43$ $0.96\pm 0.008$
AudioSeal $4.32\pm 0.29$ $0.99\pm 0.003$
AWARE $4.08\pm 0.37$ $0.97\pm 0.002$
Robustness results in Table 2 demonstrate that AWARE remains effective across diverse edits, with only small fluctuations in BER between conditions. In contrast, AudioSeal degrades sharply on spectral edits (e.g., LPF and BSF), while WavMark fails under coarse quantization and compression (PCM $8$-bit and MP3 $64$ kbps). Under pink-noise corruption, AWARE outperforms comparative methods, indicating strong resilience to background noise effects. Passage through a neural vocoder yields very low BER for AWARE, whereas baselines struggle. Temporal edits remain the hardest case. Under sample deletions and time-scale modifications AWARE incurs higher BERs, though still within a usable range.
Table 2: BER (%) under various edit/attack conditions.
Condition WavMark AudioSeal AWARE
Original 0.00 0.00 0.00
LPF 0.00 $14.58$ 0.00
HPF 0.00 $7.08$ 0.00
BSF 0.00 $33.81$ 0.95
PCM $24.46$ 1.47 1.43
MP3 $24.12$ 0.24 $0.71$
PN $28.59$ $10.89$ 1.61
RS 0.00 0.00 0.00
SD $1.43$ 0.70 $3.74$
TS $9.98$ $10.58$ 5.53
PS $50.00$ $2.55$ 0.92
NV $50.00$ $39.01$ 1.61
We again note that the results for the AWARE system are obtained without direct simulation of specific attacks during training. Robustness largely stems from the detector architecture and the BRH. Conversely, competing systems often see (and thus favor) certain distortions during training, which helps in those particular cases but can leave gaps elsewhere.
6 Ablation Studies
To evaluate effectiveness and rationale of key architectural and representational design choices, we conducted a series of ablation studies. Each experiment isolates one architectural or representational component while keeping all other settings fixed. This allows us to examine how each component, from the Bitwise Readout Head to kernel size and spectral domain selection, impacts robustness and consistency under typical audio distortions. For every ablation, we chose representative attack subsets that the ablated component is expected to affect. The results are summarized in the tables below as mean BER across those cases.
6.1 BRH vs. Fully Connected Output
In Table 3, we compare the proposed BRH with a variant that replaces it by a conventional FC output layer, which binds detection to absolute positions and thus loses time-order invariance. We focus on sample deletion and time-stretch conditions, as they most directly probe temporal robustness.
Table 3: BER (%) under deletion (SD) and time-stretch (TS) attacks for models with and without BRH.
Model SD TS
w/ BRH (proposed) 3.74 5.53
w/o BRH (FC layer) 30.91 7.84
The BRH substantially improves robustness under temporal desynchronization, as it aggregates evidence over time without depending on frame order or absolute position. This effect is particularly evident in fragment-level detection, where portions of the signal are entirely removed: the BRH can still accumulate sufficient evidence from the remaining segments to correctly infer watermark bits, while the FC variant collapses once the input continuity is broken.
6.2 Kernel Size 1 vs. 3
Next, we assess the effect of convolutional kernel size along the temporal axis. We analyze convolutional kernels of size $1$ and $3$. Kernel size $1$ avoids temporal mixing and maintains activation stability under cuts, whereas kernel size $3$ introduces context leakage across frames.
Table 4: BER (%) under temporal edits for different convolutional kernel sizes.
Kernel size SD TS
11 (proposed) 3.74 5.53
33 6.06 13.96
Results, given in Table 4, support the choice of minimal temporal context in feature extraction blocks. Restricting convolutional kernels to size $1$ keeps activations independent across frames and preserves stability under deletions and temporal warping.
6.3 With vs. Without Mel Projection
Finally, we compare the proposed Mel-based detector with a variant operating directly on STFT magnitudes. The Mel front-end performs perceptual band aggregation, which stabilizes frequency-domain statistics and aligns the representation with human-critical bands and common synthesis pipelines. Accordingly, we probe attacks that directly stress these properties: LPF/HPF, MP3, and neural vocoder resynthesis, precisely where Mel aggregation is expected to enhance the robustness.
Table 5: BER (%) under neural vocoder (NV) resynthesis and spectral/compression attacks with and without the Mel front-end.
Front-End NV LPF HPF MP3
w/ Mel (proposed) 1.61 0.00 0.0 0.71
w/o Mel (STFT only) 50.30 0.68 0.83 1.42
Results in Table 5 suggest that for spectral edits (LPF/HPF) and compression (MP3), absolute differences are modest, yet consistently favor the Mel front-end. Under NV resynthesis, the gap becomes substantial. The Mel-based detector remains highly reliable, whereas the STFT-only variant degrades sharply.
7 Conclusion
This paper advocates robustness by design for digital audio watermarking and introduces a time–order–agnostic detector with a Bitwise Readout Head. Adversarial evaluation serves as a diagnostic tool to surface right inductive biases, without relying on heavy attack simulation. The system delivers high audio quality and consistently low BER across diverse edits, often surpassing strong learning-based baselines. However, rather than pursuing “state-of-the-art” claims, particularly tenuous in this domain, given competing metrics and trade-offs, we aim to provide principled guidance on design choices tailored to the intended threat model for digital audio watermarking use cases.